Complete this table to show the pH and pOH of each solution.


(H+ is short for H3O+)


  [H+] [OH−] pH pOH

Solution A 1.0×10−5 5

Solution B 1.0×10−10 10 

Solution C 1.0×10−12

Solution D 10

Solution A [OH−]= 

M

Solution A pOH= 

Solution B [H+]= 

M

Solution B pH= 

Solution C [OH−]= 

M

Solution C pH= 

Solution C pOH= 

Solution D [H+]= 

M

Solution D [OH−]= 

M

Solution D pOH= 



1
Expert's answer
2022-04-26T12:28:02-0400

solution a pOH= 9, [OH−]=1*10^-9

solution b pH=10, [H+]=1*10^-10

solution c pH = 12, pOH = 2

solution d pOH=15, [OH−]=1*10^-15